# Enabling Physical Localization of Uncooperative Cellular Devices

### v2

@@ 5.3 Power Boosting Attack
In this evaluation, we initially execute scheduling manipulation attack before injecting TPC commands to show the feasibility of an integrated approach. Before and after the attack, we measure the transmission power of the UE using a diagnostic monitoring tool, XCAL [53], to show the increase of uplink transmission power.
@@
As a result, we confirm that while UE’s transmission power is about 10dBm before performing the attack, through the power boosting, the transmission power of UE is increased up to 23dBm (i.e., maximum transmission power). Considering the transmission power of the UE presented in Table I, it is feasible to overcome the shadow area by increasing the transmission power to 23dBm through the power boosting attack.
5.4 Demonstration in Commercial Testbed
@@
To demonstrate the practicality of scheduling manipulation and power boosting attacks, we examine the combined attack in the commercial LTE testbed. In this evaluation, we execute scheduling manipulation attack as the first step to compel the UE to generate uplink transmissions and maintain the UE’s RNTI. Subsequently, we conduct power boosting attack to increase the UE’s uplink transmission power while the scheduling manipulation attack is ongoing. In this experiment, we use three COTS devices (Table IV). As a result, we successfully conduct scheduling manipulation and power boosting attacks on all UEs in the commercial testbed. Demonstration videos illustrating the combined attack in commercial testbed are available [^2] Demo videos: S10 – https://www.youtube.com/watch?v=Z_b9eR65Q50 S20 – https://www.youtube.com/watch?v=KC2Baoi_1do ZFlip4 – https://www.youtube.com/watch?v=x02KNdxEosU&t=30s .
5.5 End-to-End Evaluation
@@
Experimental Setup.
Figure 6: Experimental setup for end-to-end evaluation.
We conduct the experiment in $15m\times 10m$ lab environment as shown in Figure 6. Note that, due to ethical concerns regarding scheduling manipulation and power boosting attacks, we built the testbed with srsRAN [47]. Even though we install srsENB using high-performance SDR, USRP X310 [51], its cellular range is inevitably limited. Therefore, we arrange the experimental setup in a spatially constrained room. We deploy two sniffers equipped with the directional antenna [28], one near the eNB and another at the border of the cellular range, respectively. As depicted in the figure, the target UE is at nine points. For each evaluation trial, we turn on extra two UEs at the same time, connecting them to the eNB. Afterward, our evaluation is designed to figure out where one of the UEs (i.e., target UE) is physically located. Especially, since srsENB doesn’t support SMS, we utilize WhatsApp for RNTI acquisition. More specifically, in this step, we generate traffic patterns without triggering any alarms on the target UE using the ‘typing notification’ feature of WhatsApp [46].
@@ Results.
TABLE V: Evaluation of end-to-end experiment.
@@
70% angluar accuracy $\sim$ ${}^{\circ}$ $\sim$ ${}^{\circ}$
@@
Table V shows the summarized performance results from the end-to-end evaluation of the various metrics. We confirm that all nine trials of end-to-end evaluation successfully pinpoint the physical location of the target UE with 70% distance error of 1.7m. Furthermore, we confirm that each UMA attack, where the four key steps are conducted in sequence, takes less than 5 minutes to determine the physical location of a single target UE from its online identity. Not surprisingly, it takes about 2 minutes to perform power boosting attack due to the process of acquiring radio configuration and manipulating subframes, as described in § 4.3.2. In multiangulation step, it also takes about 2 minutes to estimate the direction (i.e., angle) of the signal source at two sniffers and determine the location.
@@
We further analyze the effectiveness of the power boosting attack on the end-to-end evaluation. Table V presents a performance comparison between two scenarios: the end-to-end evaluation with and without the power boosting attack. The power boosting attack enhances localization performance, as evidenced by improvements in angular accuracy, localization accuracy, and arrival signal strength on the sniffers’ side. Additionally, in cases where we perform the end-to-end process without executing the power boosting attack, the physical location of the target remains undetermined at two out of nine locations, marked by gray circles in Figure 6.
@@
To investigate these differences, we measure the arrival signal strength according to the angle from the UE placed at the left gray circle in Figure 6. We rotate the antenna from ${}^{\circ}$ (i.e., pointing to the UE) to ${}^{\circ}$, using the sniffer placed at the bottom right. As depicted in Figure 7, in the absence of the power boosting attack step, the angle corresponding to the peak signal strength is not discernible, making it challenging to determine the angle of the signal source. This directly leads to errors in multiangulation, causing the attacker to pinpoint the physical location inaccurately. This observation aligns with the discussion in § 3.1, where it is noted that when the UE is in close proximity to the eNB, it generates its uplink signal with low transmission power. Additionally, Figure 7 hints that power boosting can improve the direction finding capability by reducing the noise on the sniffer side when the attacker attempts to estimate the direction of signal source. Actually, the signal-to-noise ratio (SNR) in the UE direction has been found to be 30dB with power boosting and 13dB without power boosting. A demonstration video illustrating the end-to-end evaluation is available: https://www.youtube.com/watch?v=Y0MJNQCKs7I&t=1s.
Figure 7: Impact of the power boosting on direction finding.
5.6 Power Boosting vs. Cellular Repeater
@@
We now examine the effectiveness of our strategy employing the power boosting attack to distinguish the UE’s signal from the repeater’s. Due to ethical concerns, the evaluation is conducted on the commercial testbed. This testbed comprises commercial network equipment (§ 5). The repeater’s external antenna was installed near eNB and the UE was positioned 2m away from the repeater. We then placed the sniffer 5m away from the UE and the repeater’s external antenna.
@@
We first measure the arrival signal strength at the direction of the UE and the repeater’s external antenna (Table VI) without the power boosting attack. The arrival signal strength at the direction of the UE is measured as 5dBm. At the direction of the repeater, the signal strength is equivalent to 17dBm. Next, we measure the arrival signal strength in two directions, as before, while conducting the power boosting attack. Signal strength at the direction of UE, at this time, significantly increases to 25dBm, while the arrival signal strength at the direction of the repeater remains unchanged (static). As a result, as we described in § 4.3.3, by observing an increase in arrival signal strength using a directional antenna after performing the power boosting attack, the UMA attacker can identify the UE’s true signal.
TABLE VI: Impact of cellular repeater and TPC injection.
@@
  • Distance between the repeater’s antenna and the UE is 2m
In conclusion, through the observation of an increase in arrival signal strength using a directional antenna after performing the power boosting attack, the UMA attacker can distinguish the location of the UE from that of the repeater.
5.7 Iterative Direction Finding in operational network
Lastly, we present a use case for the physical localization of uncooperative cellular devices in an operational LTE network. More specifically, similar to the scenario we described in § 3.1, we show how a localizer can find a UE in a room on the 2nd floor of a building measuring $56m\times 34m\times 15m$, camped in the urban operational network. The demostration begins by deploying a sniffer near the eNB (approximately 10m away from the tower).
To acquire the target’s RNTI, we employ SMS messages to generate traffic patterns. This is achievable by monitoring downlink traffic, as discussed in § 4.1. After acquiring the target’s RNTI, scheduling manipulation would be conducted to force the UE to generate uplink traffic and maintain the UE’s RNTI. Due to ethical considerations, however, we employ voice call service to replace the effect of scheduling manipulation and power boosting. Naturally, a voice call makes UE continuously generate uplink traffic and maintain its radio connection (i.e., fixed RTNI). Furthermore, UE’s transmission power increases in response to data throughput requirements of voice call – we empirically confirm this in § B.4 along with the specification [1]. Subsequently, we track down the physical location of the target UE by moving the sniffer’s position.
Figure 8: The process of iterative DFing and localization heatmap targeting UE inside a building.
Figure 8 illustrates the tracking process in this scenario. Initially, we estimate the direction of the signal source at the starting point (L1) and deduce the building where the target UE is suspected to be located. Then, we move the sniffer to the front of the suspected building (L2). At the second point, we investigate on which floor the target UE is located by measuring arrival signal strength, as illustrated at the bottom of Figure 8. Based on the measurement, we determine that UE is most likely on the 2nd floor. Finally, after moving the sniffer to the 2nd floor inside the building (L3), we successfully pinpoint the physical location of the target UE by tracking down the direction of signal source. During the tracking process, if the target UE’s uplink signal is undetectable or the repeater is installed, power boosting can be performed to increase the target’s transmission power.
6 Related Work
TABLE VII: Comparison with related works
ToA-based approach AoA-based approach
RNTI tracking from online identity [31]${}^{1}$ UMA
Targeted localization [31, 48] UMA
No need for H/W or RF fingerprinting [48]${}^{2}$ UMA
Generating continuous uplink signal UMA
Boosting uplink signal power UMA
Consider cellular repeater UMA
[…]
  • 1
    LTrack obtains RNTI from IMSI/TMSI, which are more challenging to acquire compared to online identities.
  • 2
    Localization performance with 78m granularity.
  • 3
    They utilize both passive and active approaches for localization.
Several studies proposed the physical localization methods against cellular devices [31, 32, 48]. These prior works and our work utilize the downlink/uplink signal for localizing the target UE. However, our work has additional capabilities in its threat model and provides novel approaches to address the challenges that have not been discussed before (refer to Table VII). Note that previous works have different goals and threat models.
ToA-based localization. Kotuliak et al[31] proposed a localization attack based on ToA measurement. Their approach exploits the unencrypted information for the transmission time alignment (i.e., TA) and the actual monitored uplink transmission time. Based on these two values, the attacker estimates the victim’s location with a 90% probability and an error of less than 6m. Tan et al[48] proposed data-plane signaling forgeries and utilized the fabricated malicious signaling to determine the position of IoT devices. Similar to the above work [31], they estimated the target’s location based on the eavesdropped TA values from multiple eNBs. However, their method, which solely relies on TA commands, shows insufficient precision (78m).
Compared to those works, our work differs in three aspects: (1) our scenario leverages the victim’s online identity, which is more easily obtainable compared to IMSI/TMSI, for RNTI tracking; (2) we introduce a scheduling manipulation attack to enforce the victim UE to transmit an uplink signal, ensuring a fixed RNTI. This approach helps overcome practical challenges during the localization attack; (3) the challenges and evaluations are built upon a real world scenario. This is important as those challenges are only identifiable in the practical environment.
AoA-based localization. The closest work to our work in the context of localization method (i.e., AoA) is LTEye [32]. The authors introduced an indoor localization attack based on AoA measurement, which utilizes a synthetic aperture radar to handle the multipath problems in indoor environment. They estimated the target UE’s location by using multiple radars at different locations. LTEye also provided an RNTI tracking method to map RNTIs assigned to the same device, which correlates the currently monitored RNTIs to the expired RNTI. However, this method necessitates an RF fingerprinting procedure with a corresponding database. While both LTEye and UMA rely on uplink signal monitoring, UMA employs a more active attacker model by exploiting the uplink transmission logic through the power boosting attack and scheduling manipulation attack. In contrast, LTEeye solely depends on passive monitoring of uplink traffic. Note that, the two proposed novel attacks enable more stable location tracking for both LTEye and UMA. In addition, UMA is designed to targeted attacks, but LTEye cannot be utilized in targeted localization attacks (i.e., estimating the location of a specific victim) since it cannot identify a specific device.
Other approaches. Lakshmanan et al[34] demonstrated unprivileged location tracking by exploiting Carrier Aggregation (CA) side channels in cellular networks. However, it requires collecting CA measurements in all possible target paths, limiting the capability of localizers when attempting to trace cellular devices. Besides cellular localization, Abedi et al[7] presented uncooperative WiFi localization through the injection of fake packets. Another example, Blue’s Clues [50] introduced a method to track Bluetooth devices even when they are in non-discoverable mode. In addition, Zhang et al[57] presented the physical localization of cellular devices by impersonating the victim and replaying messages to the victim’s Bluetooth device. Nonetheless, since these methods exploit short-range communication protocols, the attacker’s ability to trace the victim is limited to a constrained range.
7 Discussion
Novelty and goal. Our main contribution is identifying and addressing realistic challenges associated with physically tracking uncooperative devices in cellular networks through two novel approaches; scheduling manipulation and power boosting. While various localization techniques for precise tracking have been actively studied, our perspective on achieving reliable uncooperative cellular localization has not been previously discussed. We have demonstrated the feasibility of our approaches through evaluations in three environments, within legal boundaries. Based on our findings and evaluations, GSMA has acknowledged our contributions and confirmed that they are exploitable for localization.
While we are not primarily focusing on introducing new localization techniques or enhancing performance, our approach presents an opportunity for performance improvement. In this work, although we demonstrated the effectiveness of UMA against the received signal strength-based localization, our approach is also applicable to Channel State Information (CSI) based localization methods [40, 56, 16], as well as other physical-layer techniques [31, 32]. In such integrated scenarios, the performance of localization may be enhanced by mitigating noise and interference through scheduling manipulation and power boosting, as more signals with higher signal gains tend to result in better channel quality.
Although BSR exploitation, a key component of the scheduling manipulation attack, has been discussed in [48], our approach differs in terms of attack procedures and goals. In [48], the attacker attempts to achieve uplink resource draining by sending forged BSR repeatedly, as a single-protocol attack. In our approach, the attacker impersonates the victim UE and maliciously injects both scheduling requests and BSR of the victim to achieve two objectives necessary for uncooperative physical localization: 1) compel the victim to continuously generate uplink traffic and 2) maintain the victim’s RNTI.
UMA with prior studies. LTEye [32] is a representative study employing AoA for cellular localization, similar to UMA. Thus, our approaches are easily applicable to LTEye. As an initial step of targeted localization with UMA and LTEye, the attacker acquires the target UE’s RNTI associated with its online identity by generating a silent traffic pattern. Then, the attacker executes the scheduling manipulation attack to force the UE to continuously generate uplink traffic and maintain its RNTI. This enables LTEye to effectively perform uncooperative physical localization without its practical requirements; 1) the target UE should be in an active status while the SAR’s mobile antenna rotates 360°and 2) the RNTI mapping process with RF fingerprinting is needed when the target’s RNTI is reassigned. When the attacker faces challenges stemming from low uplink transmission power or the presence of the repeater, the attacker performs the power boosting attack to improve localization performance.

### v3

@@ 5.3. Power Boosting Attack
We investigate whether the TPC injection indeed increases UE’s transmission power up to the maximum level. To demonstrate the effectiveness of the power boosting attack in UMA, we run the experiment in the testbed environment, using six COTS devices. In this evaluation, we initially execute the scheduling manipulation attack before injecting TPC commands to show the feasibility of an integrated approach. Before and after the attack, we measure the transmission power of the UE using a diagnostic monitoring tool, XCAL (Accuver, 2015), to show the increase of uplink transmission power.
@@
As a result, we confirm that while UE’s transmission power is about 10dBm before performing the attack, through the power boosting, the transmission power of UE is increased up to 23dBm (i.e., maximum transmission power). Considering the transmission power of the UE presented in Table 1, it is feasible to overcome the shadow area by increasing the transmission power to 23dBm through power boosting.
Figure 5. Experimental setup for e2e evaluation.
5.4. Demonstration in Commercial Testbed
@@
We now move to the commercial LTE testbed and examine the feasibility of the combined attack (scheduling manipulation and power boosting attack). We execute the scheduling manipulation attack as the first step to compel the UE to generate uplink transmissions and maintain the UE’s RNTI. Subsequently, we conduct the power boosting attack to increase the UE’s uplink transmission power while scheduling manipulation is ongoing. As a result, we confirm that the combined attack works for three types of devices and commercial network equipment. Demonstration videos illustrating the combined attack are available on our website (UMA, 2024).
5.5. End-to-End Evaluation
@@
Setup. We conduct the experiment in $15m\times 10m$ lab environment as shown in Figure 5. Note that, due to ethical concerns regarding scheduling manipulation and power boosting attacks, we built the testbed with srsRAN (Systems, 2019). Even though we install srsENB using high-performance SDR, USRP X310 (Research, 2014), its cellular range is inevitably limited. Therefore, we arrange the setup in a spatially constrained room. We deploy two sniffers equipped with the directional antenna (Aaronia, 2023), one near the eNB and another at the border of the cellular range, respectively. As depicted in the figure, the target UE is at nine points. For each trial, we turn on extra two UEs at the same time, connecting them to the eNB. Afterward, our evaluation is designed to figure out where one of the UEs (i.e., target UE) is physically located. Especially, since srsENB doesn’t support SMS, we utilize WhatsApp for RNTI acquisition (Shaik et al., 2016).
@@ Results.
Table 5. Evaluation of end-to-end experiment.
@@
70% angluar accuracy $\sim$ 16 $\sim$ 12
@@
Table 5 shows the summarized performance results from the end-to-end evaluation across various metrics. We confirm that all nine trials of end-to-end evaluation successfully pinpoint the physical location of the target UE with 70% distance error of 1.7m. Furthermore, we confirm that each UMA attack, where the four key steps are conducted in sequence, takes less than 5 minutes to determine the physical location of a single target UE from its online identity. Not surprisingly, it takes about 2 minutes to perform the power boosting attack due to the process of acquiring radio configuration and manipulating subframes, as described in § 4.3.2. In multiangulation step, it also takes about 2 minutes to estimate the direction (i.e., angle) of the signal source at two sniffers and determine the location.
@@
We further analyze the effectiveness of the power boosting attack on the end-to-end evaluation. Table 5 presents a performance comparison between two scenarios: the end-to-end evaluation with and without the power boosting attack. The power boosting attack enhances localization performance, as evidenced by improvements in angular accuracy, localization accuracy, and arrival signal strength on the sniffers’ side. Additionally, in cases where we perform the end-to-end process without executing the power boosting attack, the physical location of the target remains undetermined at two out of nine locations, marked by gray circles in Figure 5.
Figure 6. Impact of power boosting on DF-ing.
@@
To investigate these differences, we measure the arrival signal strength according to the angle from the UE placed at the left gray circle in Figure 5. We rotate the antenna from 0 (i.e., pointing to the UE) to $\pm$50°, using the sniffer placed at the bottom right. As depicted in Figure 6, in the absence of the power boosting attack step, the angle corresponding to the peak signal strength is not discernible. This directly leads to errors in determining the angle of the signal source, causing the attacker to pinpoint the physical location inaccurately. This observation aligns with the discussion in § 3.1, where it is noted that when the UE is in close proximity to the eNB, it generates its uplink signal with low transmission power. Additionally, Figure 6 hints that power boosting can improve the direction finding capability by reducing the noise on the sniffer side when the attacker attempts to estimate the direction of signal source. Actually, the signal-to-noise ratio (SNR) in the UE direction has been found to be 30dB with power boosting and 13dB without power boosting.
5.6. Power Boosting vs. Cellular Repeater
@@
We now examine the effectiveness of our strategy employing the power boosting attack to distinguish the UE’s signal from the repeater’s. Due to ethical concerns, the evaluation is conducted on the commercial testbed. The repeater’s external antenna was installed near eNB and the UE was positioned 2m away from the repeater. We then placed the sniffer 5m away from the UE and the repeater’s external antenna.
@@
We first measure the arrival signal strength at the direction of the UE and the repeater’s external antenna (Table 6) without the power boosting attack. The arrival signal strength at the direction of the UE is measured as 5dBm. At the direction of the repeater, the signal strength is equivalent to 17dBm. Next, we measure the arrival signal strength in two directions, as before, while conducting the power boosting attack. Signal strength at the direction of UE, at this time, significantly increases to 25dBm, while the arrival signal strength at the direction of the repeater remains unchanged (static). As a result, by observing an increase in arrival signal strength after performing the power boosting attack, the UMA attacker can identify the UE’s actual signal.
Table 6. Impact of cellular repeater and TPC injection.
@@
  • Distance between the repeater’s antenna and the UE is 2m.
5.7. Attack Feasibility
Validity of experimental setup. Due to legal and ethical constraints, we were unable to conduct end-to-end evaluations of UMA, including scheduling manipulation and power boosting attacks, in operational networks. These attacks could affect operational networks and normal user devices since they require signal injection over both downlink and uplink channels.
To address this limitation, we designed our experimental setup to mirror real-world settings as closely as possible. To demonstrate the attack feasibility, we used combinations of COTS devices and eNBs for various experiments, as shown in § 5: 1) Nine COTS devices with different baseband chipsets manufactured by five vendors covering over 90% of the market share (Analytics, 2022), and 2) A commercial testbed featuring an eNB from Nokia, a leading eNB vendor, and a lab testbed built with open-source srsRAN (Systems, 2019) supporting a full-stack eNB.
Standard-compliant feature exploitation. Our approaches exploit only standard-compliant features in LTE, including scheduling request, BSR, and TPC. Therefore, they should be effective on any standard-compliant eNBs and UEs (i.e., universal cellular environments). Additionally, GSMA stated, “We acknowledge that the mechanisms of power boosting attack and scheduling manipulation attack can make localization attacks more reliable,” upon our responsible disclosure. This remark confirms that the observed result aligns with the expected standard-compliant behavior.
Power boosting with SigOver. Power boosting leverages SigOver (Yang et al., 2019), a well-known signal overshadowing attack shown to be effective in commercial networks (Erni et al., 2022; Yang et al., 2019; Tan et al., 2021; Kotuliak et al., 2022). It injects malicious signals that overshadow legitimate ones, making the UE receive only the injected malicious signals due to capture effect and synchronization with targeted legitimate signals. Exploiting the capture effect and synchronization method of SigOver, a UMA attacker can increase the transmission power of the victim UE to the maximum.
6. Discussion
Other potential attacks using scheduling manipulation. Scheduling manipulation induces continuous uplink traffic generation on the target UE, potentially accelerating battery drain. To quantify this impact, we measured electric current levels in the UE’s circuitry under both attack and normal conditions using a battery monitoring application (Paget96 Projects, 2019). Electric current directly correlates with battery consumption. During the scheduling manipulation attack, the average electric current level was measured at 340mAh, compared to 250mAh during normal cellular connection. This significant increase in the current level demonstrates that the scheduling manipulation attack substantially accelerates UE battery drain by forcing continuous uplink traffic generation.
We also analyzed the feasibility of the overcharging attack. Our analysis confirmed that scheduling manipulation cannot affect billing (i.e., charging) since the target UE’s traffic generated by the attack consists only of padding, not actual data payload. In cellular networks, charges are incurred, when the Service Data Unit (SDU), the data payload for cellular service usage, is included in the packet and forwarded to the Packet Data Network Gateway (P-GW), specifically the Policy and Charging Enforcement Function (PCEF) in the core network, after decoding and decryption by the eNB. However, packets generated by scheduling manipulation don’t include an SDU (refer to our website (UMA, 2024)). Moreover, the attacker’s packets aren’t forwarded to the P-GW, as the attacker lacks the target’s key for proper packet encryption.
Lastly, scheduling manipulation could potentially lead to the uplink resource draining attack, as demonstrated in (Tan et al., 2021), where spoofing BSR results in unfair resource assignment.
Limitations. First, the entire UMA was not tested end-to-end in operational cellular networks due to ethical concerns. Practical issues during physical localization may arise, such as multipath effects, especially in urban environments in reality. Fortunately, existing research has explored strategies for radio direction detection (Treichler and Agee, 1983; Vanderveen et al., 1998) and indoor localization (Chintalapudi et al., 2010; Otsason et al., 2005; Zafari et al., 2019; Kumar et al., 2014). We believe integrating those algorithms can enhance localization accuracy in such environments. Second, while UMA should work in 5G networks based on specifications, it requires developing the necessary toolkit: a) 5G sniffer: LTESniffer (Hoang et al., 2023) supports high-performance downlink and uplink sniffing in LTE, but existing 5G sniffer (Ludant et al., 2023) lacks uplink channel monitoring capabilities; b) 5G signal overshadowing: This has not been implemented yet; c) New 5G features: Wider bandwidth, beamforming, and massive MIMO may affect UMA design for 5G. Despite the requirements, attackers can still conduct UMA by downgrading the UE’s connection through bidding down attacks. Finally, UMA could potentially be detected by the victim’s device (§ 7.1), allowing them to turn off their device to avoid tracking. However, unless eNB/UE vendors implement attack detection, victims would need specialized equipment such as a sniffer, diagnostic monitoring tool (e.g., XCAL), and algorithms to differentiate attack packets from legitimate ones.
Other application. UMA can be used in rescue operations for individuals inevitably unable to cooperate with localization. People who are injured, kidnapped, or cognitively impaired, may be unable to request rescue or report their location.
7. Countermeasure
7.1. Mitigation for Scheduling Manipulation
Mitigation at baseband. When the UE receives DCI 0 messages containing uplink resource allocation, it transmits packets over the uplink channel even if it did not request such allocation. To mitigate this issue, baseband vendors can implement a condition where data is transmitted only if the device previously requested uplink grants or transmitted a scheduling request. Alternatively, the device simply doesn’t transmit data if there is no pending data in its buffer.
Detection at base station. During scheduling manipulation, two anomalous behaviors occur; 1) The target UE reports it has no data to send in the current buffer, yet it repeatedly transmits packets over the uplink channel. 2) The uplink traffic contains only padding data as payload. These behaviors could potentially be detected at the eNB as anomalous.
Reallocation of radio config. The eNB can enhance security by reallocating user-dedicated radio configurations over a secure channel after completing the security mode procedure. This reallocation can be done through RRC Connection Reconfiguration message. By using this encrypted message to reassign the radio configuration, it becomes more challenging for an attacker to identify the target UE’s specific radio configuration which is crucial for scheduling manipulation.
7.2. Long-Term Mitigation
Long-term mitigation requires modification of specification. Fundamentally, UMA exploits three vulnerabilities in lower-layer protocols of current cellular networks; 1) Lack of confidentiality in DCI messages, allowing location tracking, 2) lack of integrity protection in DCI messages, enabling the power boosting attack, and 3) scheduling manipulation attack due to lack of integrity protection in the uplink scheduling process. As 3) could be defended with short-term mitigation (§ 7.1), we focus on 1) and 2) here. Specifically, we discuss confidentiality and integrity protection for DCI 0, which manages both uplink resource allocation and transmission power.
Security protections for lower-layer cellular protocols have been explored (Tan et al., 2021; Yang et al., 2019), but significant challenges remain in protecting DCI messages; 1) DCI is a PHY layer message, typically neither encrypted nor integrity protected by design, 2) only restricted space is reserved for DCI 0 messages (maximum 37 bits: up to 13 bits for resource allocation and 2 bits for TPC) to save radio resources, making it difficult to add message authentication code or digital signatures for integrity protection, 3) DCI is fundamental to cellular network operations, managing radio resources and controlling various communication aspects. Thus, protecting DCI messages would likely affect a significant portion of cellular technology design. Considering these challenges, we leave developing specific protection mechanisms for future work.
8. Related Work
Table 7. Comparison with existing models.
Steps
Existing Physical
localization (Kotuliak et al., 2022; Kumar et al., 2014)
UMA
1) RNTI acquisition
From IMSI (Kotuliak et al., 2022),
RF fingerprinting (Kumar et al., 2014)
From online identity
((Hong et al., 2018; Hussain et al., 2019; Kohls et al., 2019; Kune et al., 2012; Ludant et al., 2023; Rupprecht et al., 2019; Shaik et al., 2016))
2) Location tracking
ToA (Kotuliak et al., 2022), AoA (Kumar et al., 2014)
measurements
AoA measurements
2-1) Fixing RNTI
C1: Limited tracking time
(RNTI duration 15$\sim$30 sec)
Unlimited tracking time
2-2) Forcing uplink traffic
C1: Untraceable during
no traffic
Traceable, regardless of
UE’s status
2-3) Power boosting C2: Limited detection range
Improved signal detection
range (70$\uparrow$) and SNR
3) Repeaters
C3: Tracing repeaters,
instead of UE
Distinguishing UE and
repeater
  • $\ast$ Applicable to ToA measurements.
  • $\dagger$ Scheduling manipulation.
  • $\ddagger$ Refer to the shadow area graph on our website (UMA, 2024).
Physical localization. Kotuliak et al(Kotuliak et al., 2022) proposed a localization attack based on ToA measurement. Their approach exploits unencrypted information for the transmission time alignment (i.e., TA) and the actual monitored uplink arrival time. This ToA-based localization requires an additional process for fingerprinting the target device to account for hardware errors. The closest work to ours in the context of localization method (i.e., AoA) is LTEye (Kumar et al., 2014). It introduced indoor localization, utilizing a SAR to handle multipath issues.
Compared to UMA, existing models on physical localization didn’t consider realistic scenarios as described in Table 7. First, although they also provided RNTI acquisition for physical localization, LTrack (Kotuliak et al., 2022) leverages the target’s IMSI/TMSI, which is more difficult to obtain compared to online identity in realistic scenarios, and LTEye (Kumar et al., 2014) necessitates an RF fingerprinting procedure with a corresponding database to trace the RNTIs. Second, they have limited opportunities for physical localization. They can trace the target UE only while it is in active status, connected to the eNB. This duration can be only 15–30 seconds, RNTI duration, in the worst case, restricting the attacker’s localization capabilities. Additionally, even if the target UE is connected to the eNB, it may not transmit uplink traffic if there is no data to send in its current buffer, making the attacker unable to trace the target UE. Finally, they cannot correctly measure ToA or AoA when the target UE transmits its signals with low power or repeaters relay the target’s signals. This restricts the attacker’s signal detection range and makes physical localization difficult.
To overcome these challenges in the physical localization of uncooperative cellular devices, they should adopt our approach. Thanks to UMA, attackers can effectively achieve reliable fine-grained localization in realistic scenarios.